Using the mathematical model: to calculate velocity of a car accelerating from rest in straight line. Equation is v(t) = A (1-e ^ -t/t max speed) the ^ means to the power of. v(t) is the instantaneous velocity of the car (m/s), t is the time in seconds, tmaxspeed is the time to reach the maximum speed in seconds and A is a constant. Find out how to: 1. Identify the units of the coefficient A, 2. physical meaning of A, 3. velocity of the car at t=0 and 4. asymptote of this function as t ⮕ ∞?
1
Expert's answer
2020-01-10T09:18:31-0500

"v(t)=A(1-e^{-t\/t_o})" is the equation given for velocity, where;

"t_o" = time to reach the maximum speed.


1.Except A all other terms in the right hand side of the equation are dimensionless constants.

Hence, A must be of the dimensions same as that of velocity.

"\implies" A has unit same as velocity which is "(m\/s)" .


2."lim_{v \to \infty} v(t)=A(1-e^{-\infty})=A(1-0)=A" Thus physical significance of A is that it is the terminal velocity(as t tends to infinity, velocity tends to A)


3. "v(t)=A(1-e^{-t\/t_o})"

"\implies v(0)=A(1-1)=0m/s"


4. As t tends to infinity,v(t) is same as terminal velocity that is equal to A.

Thus, asymptote of this function as "t \to ∞" is "v=A" .